Find the area of the parallelogram with vertices A(-3, 3), B(-1,
If a = (2, -1, 4) and b = (7, 2, 1), find the following. a xb = b x a =
Find the cross product a x b. a = i+ 2j - 4k, b = -i + 5k
Answer #1

Question 1 is incomplete. Only one vertex is given. So I have solved the remaining questions.sol A = (2,-1,4) and b =( 7, 2, 1) from the definition of cross product we have a no axb = (as b3 - 83 b) î+ (azb, - aiba 9 +ash+29-4% , b= 2 +5% Writing ng the above given rector in co-oredinate as a = 21,2,-4) 2 ball; 0,5%. Now, s 3 from the defini